OpenStudy (anonymous):

4/5x- 3 3/4x= 2 1/2 SOLVE

OpenStudy (mimi_x3):

is it\[\large \frac{4}{5} x-3\frac{3}{4x} = 2\frac{1}{2} \]?

OpenStudy (anonymous):

That's not in the options?

OpenStudy (mimi_x3):

lol, no is that the question?

OpenStudy (anonymous):

A. -1 9/50

OpenStudy (anonymous):